If X and Y are independent and identically distributed uniform random variables on (0,1) compute the joint density of

U = X+Y, V = X/(X+Y)

Part A,

The state space of (U,V) i.e. the domain D over which fU,Y (u,v) is non-zero can be expressed as

(D = {(u,v) \epsilon R x R] 0 < h1(u,v) < 1, 0 < h2(u,v) < 1} where x = h1 (u,v) and y = h2 (u,v)

Find h1(u,v) = (write a function in terms of u and v)

Find h2(1,0.25)

Part B,

For (u,v) \epsilon D, fU,V(u,v) = (answer)
Answer #1

We are applying transformation g: (0,1)2 + R2 such that g(x,y) = (u, v) = (2+y, +y Using the theorem about the transformation

uv = x
h1(u,v) = uv
y = U - x = U - UV = U(1-V)

hence
h2(u,v) = U(1-V)